OpenStudy (anonymous):

A baseball team plays in a stadium that holds 51,000 spectators. With ticket prices at $10, the average attendance had been 38,000. When ticket prices were lowered to $8, the average attendance rose to 42,000. Find the demand function (price p as a function of attendance x), assuming it to be linear??

OpenStudy (agreene):

take the information as a pair of points, then write the equation. (p,x): (10,38000) , (8,42000) I'd suggest using point slope form because it's easier.

OpenStudy (anonymous):

i tried that and got -4000x+78000 and it was wrong

OpenStudy (agreene):

\[y-y_1=m(x-x_1)\] \[y-8=\frac{\Delta y}{\Delta x}(x-42000)\] \[y=\frac{10-8}{38000-42000}(x-42000)+8\] \[y=\frac{58000-x}{2000}\]

OpenStudy (agreene):

thats what i got anyway